# NoMIRACL: Knowing When You Don’t Know for Robust Multilingual Retrieval-Augmented Generation

### v2

@@ Abstract
Retrieval-augmented generation (RAG) grounds large language model (LLM) output by leveraging external knowledge sources to reduce factual hallucinations. However, prior works lack a comprehensive evaluation of different language families, making it challenging to evaluate LLM robustness against errors in external retrieved knowledge. To overcome this, we establish NoMIRACL, [^1] NoMIRACL: https://huggingface.co/datasets/miracl/nomiracl a human-annotated dataset for evaluating LLM robustness in RAG across 18 typologically diverse languages. NoMIRACL includes both a non-relevant and a relevant subset. Queries in the non-relevant subset contain passages judged as non-relevant, whereas queries in the relevant subset include at least a single judged relevant passage. We measure LLM robustness using two metrics: (i) hallucination rate, measuring model tendency to hallucinate an answer, when the answer is not present in passages in the non-relevant subset, and (ii) error rate, measuring model inaccuracy to recognize relevant passages in the relevant subset. In our work, we measure robustness for a wide variety of multilingual-focused LLMs and observe that most of the models struggle to balance the two capacities. Models such as LLAMA-2, Orca-2, and FLAN-T5 observe more than an 88% hallucination rate on the non-relevant subset, whereas, Mistral overall hallucinates less, but can achieve up to a 74.9% error rate on the relevant subset. Overall, GPT-4 is observed to provide the best tradeoff on both subsets, highlighting future work necessary to improve LLM robustness. [^2] Baselines and model evaluation code are available at: https://github.com/project-miracl/nomiracl
@@ 1 Introduction
Retrieval Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021; Borgeaud et al., 2022) is a promising way to incorporate external knowledge via a first-stage retrieval system. RAG instills information from reliable knowledge corpora (provided as external passages) to generate accurate and faithful responses Shuster et al. (2021); Gao et al. (2023b); Liu et al. (2023). Ever since the advent of large language models (LLMs), such as GPT-3 Brown et al. (2020) or LLAMA-2 Touvron et al. (2023), they are the de-facto choice for generation in RAG, due to their unprecedented progress in text generation and understanding Brown et al. (2020); Li et al. (2021); Chang et al. (2023); Guo et al. (2023). RAG grounds the LLM-generated output, thereby avoiding previously observed factual hallucination Maynez et al. (2020); Raunak et al. (2021) and outdated knowledge Cao et al. (2021); He et al. (2023) in LLMs.
@@
A challenging issue in RAG is to provide robust and reliable LLM-generated output. The generation stage is dependent on the first-stage information retrieval system. The retrieval system poses challenges in accurately retrieving relevant information when evaluated either on zero-shot domains Thakur et al. (2021) or low-resource languages Zhang et al. (2023). The incorrect or non-relevant information contained in retrieved passages can frequently mislead the LLM to hallucinate Adlakha et al. (2023); Chen et al. (2023); Shi et al. (2023); Yoran et al. (2023); Yu et al. (2023). Prior work lacks a comprehensive evaluation of LLM reasoning capabilities in multiple languages. As a result, it still remains unclear to which extent LLMs hallucinate across both high- or low-resource languages.
To facilitate research in this direction, we present NoMIRACL, a large multilingual human-annotated dataset containing over 56,000 (including both non-relevant and relevant samples) to evaluate LLM robustness against errors in first-stage external information, i.e., retrieved passages, across 18 typologically diverse languages. To construct the dataset, we hired 31 native speakers as human annotators Zhang et al. (2023). NoMIRACL contains two subsets, non-relevant and relevant. The non-relevant subset contains all queries with no-known answers, i.e., all top-$k$ retrieved passages manually judged as non-relevant. Conversely, the relevant subset contains queries with known answers, i.e., at least one of the top-$k$ passages is manually judged as relevant.
To better understand the LLM robustness in NoMIRACL, we conduct experiments with several existing powerful and multilingual-focused LLMs (e.g., GPT-4, Mistral, LLAMA-2). We conduct our experiments using the top-$k$ oracle passages retrieved using a hybrid retrieval system from a language-specific Wikipedia corpus Zhang et al. (2023). We use a zero-shot “vanilla” prompt template for prompting all LLMs. Our key findings are as follows: First, a majority of LLMs such as LLAMA-2, Orca-2, and FLAN-T5 observe a suprisingly high 88% hallucination rate on the non-relevant subset. Second, Mistral deals with hallucination well but performs poorly on the relevant subset. Overall, GPT-4 is found to provide the optimal performance tradeoff across both subsets.
Additionally, we conduct an empirical analysis on the complete NoMIRACL English to analyze the blind spots of LLMs, where we observe LLAMA-2-7B and 13B are interestingly found to repeat the query and prompt instructions on average by at least 25%. Mistral-7B always provides a rationale in their output generation by over 88%. This empirical analysis can be used to discover hallucination and error patterns and develop effective methods to alleviate them in LLMs.
@@ 2 Background and Problem Identification
RAG Background. Retrieval-augmented generation (RAG) (Lewis et al., 2020; Guu et al., 2020) involves a two-stage inference pipeline. In the first stage, given the retrieval system and the user query, the retrieval system provides the subset of top-$k$ passages retrieved from an external data corpus $C$. For the next stage, the user query along with the top-$k$ retrieved passages are provided to the LLM which generates an output summarizing the answer for the query, citing relevant passages in its output.
@@ 2.1 Robustness Evaluation
Definitions. NoMIRACL contains two subsets, we denote them as either non-relevant ($\mathrm{F}$) or relevant ($\mathrm{T}$). The non-relevant subset contains queries with no-known answers, i.e., all top-$k$ passages are non-relevant, whereas the relevant subset contains queries with known answers, i.e., at least one of the top-$k$ passages is relevant. Similarly, we denote the LLM prediction as either positive ($\mathrm{P}$) indicating the model finds the passage relevant to answer the query and similarly negative ($\mathrm{N}$) denotes the model does not find any passage relevant (i.e., containing the answer) for the query.
@@ 3.1 Data Construction Procedure
NoMIRACL, a supplementary dataset, was created in the course of developing MIRACL Zhang et al. (2023). NoMIRACL data construction occured in two stages, following Zhang et al. (2023); Clark et al. (2020). In the first stage, the annotator (a native language speaker) is asked to generate a well-formed query for each individual prompt text. Each prompt is a short text snippet containing the first 100 words from a language-specific Wikipedia corpus. Next, for each human-generated query, top-$k$ passages are retrieved from the Wikipedia corpus using a hybrid multilingual retrieval system (more details in §4.1). In the second stage, annotators assess the relevance of the top-$k$ query–passage pairs as a binary judgment, either relevant (relevance = $1$) or non-relevant (relevance = 00). For additional details in data construction, such as quality control, we would like the refer the reader to Appendix C.
@@
Relevant subset. Queries with known answers, i.e. at least one of top-$k$ retrieved passages is marked by the annotator as relevant to provide sufficient information to answer the query. We construct the NoMIRACL relevant subset with queries with at least one relevant passage, i.e., all query-passage pairs have been judged as either relevant with a relevance of $1$ or non-relevant with relevance of 00.
@@ 3.3 Languages Covered and Dataset Usage
NoMIRACL covers 18 diverse typological languages Zhang et al. (2023). We cover 10 language families (from Niger-Congo to Indo-European), and 11 scripts (from Latin to Devanagari) covering diversity from the perspective of linguistic characteristics. The languages along with their ISO codes are listed below: Arabic (ar), Bengali (bn), German (de), English (en), Spanish (es), Persian (fa), Finnish (fi), French (fr), Hindi (hi), Indonesian (id), Japanese (ja), Korean (ko), Russian (ru), Swahili (sw), Thai (th), Yoruba (yo), Chinese (zh).
From Table 1, we observe an uneven amount of queries present in each language. To avoid this non-uniformity and due to additional budget constraints (see subsection 4.3), in our experiments, we limit the maximum number of 250 queries for each language (if available) and subset in NoMIRACL.
@@ 4.1 Evaluation Setup and Metrics
In NoMIRACL, we evaluate LLM hallucination or error, using an input query, a vanilla prompting technique and top-$k$ (oracle) retrieved passages.
@@
Evaluation Objective. In our work, we evaluate LLM robustness as a response string $y$ in a binary classification task. Following prior evaluation strategies in Adlakha et al. (2023); Yu et al. (2023), we use the input query $q_{i}$, a vanilla prompt template $Q$, and a set of top-$k$ annotated passages $P_{k}$. We prompt the LLM to evaluate whether $q_{i}$ can be answered using any passage in $P_{k}$. The LLM generates a binary output containing either $y$ = “I don’t know” as $\mathrm{N}$ or “Yes, answer is present” as $\mathrm{P}$. The output is tagged “Invalid” if it does not fall in either one of the above. Recall from (§2.1), we calculate the hallucination rate (in %) = $\mathrm{FP}/(\mathrm{FP}+\mathrm{TN})$, which measures the error in rejecting information from non-relevant passages and the error rate (in %) = $\mathrm{FN}/(\mathrm{FN}+\mathrm{TP}$), which measures the error in identifying relevant passages amongst noisy ones.
@@ 4.2 Evaluation Models
We evaluate eleven state-of-the-art LLMs with a strong focus of multilingual instruction capabilities, including both open and closed-sourced. All model checkpoints can be found in Table 3.
(1) OpenAI: We include both closed-book LLM variants: GPT-3.5-turbo and GPT-4 OpenAI (2023) using the Azure OpenAI service. (2) Mistral: We include two variants: (i) Mistral-7B-Instruct-v0.2, a 7B instruction-tuned parameter model Jiang et al. (2023) and (ii) Mixtral-8x7B-v0.1, a sparse ixture-of-Expert (MoE) model Jiang et al. (2024). (3) Orca-2: In the Orca-2 series Mitra et al. (2023), we include both Orca-2-7B and Orca-2-13B. (4) Aya: Aya-101 Üstün et al. (2024) is a recently introduced multilingual LLM containing 13B parameters and trained with 101 languages. (5) LLAMA-2: In the LLAMA-2 series Touvron et al. (2023), we include three chat variants: Llama-2-7b-chat-hf, Llama-2-13b-chat-hf and Llama-2-70b-chat-hf instruction tuned chat models. (6) FLAN-T5: One of the earliest instruction-tuned LMs, we include flan-T5-xxl containing 11B parameters Chung et al. (2022).
@@ 4.3 Experimental Settings
We execute the generation of GPT-4 and GPT-3.5-turbo, using the OpenAI service (API version 2023-05-15) deployed on Microsoft Azure. [^5] We compared Azure API with OpenAI API across four languages in NoMIRACL and observed no noticeable difference between different GPT-4 API version providers. For a fair comparison across all LLMs, we maintain a maximum input sequence length of 4096 tokens. We set a low-temperature score = 0.1 for a deterministic LLM output, and a top-p sampling ratio = 0.95. We output a maximum of $50$ tokens.
I will give you a question and several contexts containing information about the question. Read the contexts carefully. If any of the contexts answers the question, respond as either ‘‘Yes, answer is present’’ or ‘‘I don’t know’’:
@@
Figure 5: Hallucination rate (in %) = $\mathrm{FP}/(\mathrm{FP}+\mathrm{TN}$) on the non-relevant subset in NoMIRACL test split. The non-relevant subset contains queries with no known answers, i.e., all top-$k$ (where $k=10$) passages are judged by a human annotator as non-relevant. A majority of LLMs (except Mistral) hallucinate on the non-relevant subset. Lower the hallucination rate is better. The best model in each category is plotted (see Figure 8 for all models).
@@
Vanilla Prompting. The choice of prompt significantly influences the performance and LLMs have been shown brittle to prompting variations, training examples or long context setups Liu et al. (2023). In this work, we evaluate all baselines using a zero-shot monolingual listwise prompting strategy. We construct a vanilla prompt template using all top-$k$ (oracle) passages (available in NoMIRACL) as a list of contexts along with the input query, both in the same language. We evaluate all LLMs zero-shot, as we cannot fit few-shot exemplars due to insufficient context length (maximum of 4096 sequence length). Our template provides a short description in English describing the task (Ahuja et al., 2023). Our vanilla prompt template used in our experiments is shown in Figure 4.
Reducing Costs. Running GPT-4 is expensive and LLAMA-70B is rather slow at inference. For long contexts and low-resource languages, the costs even can multiply. To limit our experimental cost, we did not exceed our prompt above 4096 tokens. To effectively fit all $k=10$ passages within the vanilla zero-shot prompt, we truncate each passage to use the first 375 tokens. Next, due to budget constraints, we keep our evaluation to a maximum of 250 randomly sampled queries for all languages in both NoMIRACL relevant and non-relevant split. In our experiments, we ended up producing $\approx$20K API calls producing an expense of \$1,474 (in USD) including miscellaneous costs.

### v3

# Knowing When You Don’t Know”: A Multilingual Relevance Assessment Dataset for Robust Retrieval-Augmented Generation

@@ Abstract
Retrieval-Augmented Generation (RAG) grounds Large Language Model (LLM) output by leveraging external knowledge sources to reduce factual hallucinations. However, prior work lacks a comprehensive evaluation of different language families, making it challenging to evaluate LLM robustness against errors in external retrieved knowledge. To overcome this, we establish NoMIRACL, a human-annotated dataset for evaluating LLM robustness in RAG across 18 typologically diverse languages. NoMIRACL includes both a non-relevant and a relevant subset. Queries in the non-relevant subset contain passages judged as non-relevant, whereas queries in the relevant subset include at least a single judged relevant passage. We measure relevance assessment using: (i) hallucination rate, measuring model tendency to hallucinate, when the answer is not present in passages in the non-relevant subset, and (ii) error rate, measuring model inaccuracy to recognize relevant passages in the relevant subset. In our work, we observe that most models struggle to balance the two capacities. Models such as LLAMA-2 and Orca-2 achieve over 88% hallucination rate on the non-relevant subset. Mistral and LLAMA-3 hallucinate less but can achieve up to a 74.9% error rate on the relevant subset. Overall, GPT-4 is observed to provide the best tradeoff on both subsets, highlighting future work necessary to improve LLM robustness. NoMIRACL dataset and evaluation code are available at: https://github.com/project-miracl/nomiracl.
Knowing When You Don’t Know”: A Multilingual Relevance Assessment Dataset for Robust Retrieval-Augmented Generation
Nandan Thakur1, Luiz Bonifacio1,3, Xinyu Zhang1, Odunayo Ogundepo1, Ehsan Kamalloo1, David Alfonso-Hermelo2, Xiaoguang Li2, Qun Liu2, Boxing Chen2, Mehdi Rezagholizadeh2, Jimmy Lin1 1 David R. Cheriton School of Computer Science, University of Waterloo, Canada 2 Huawei Noah’s Ark Lab   3 FEEC-Unicamp, Brazil
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021; Borgeaud et al., 2022) is a promising way to incorporate external knowledge via a first-stage retrieval system. RAG instills information from reliable knowledge corpora (provided as external passages) to generate accurate and faithful responses Shuster et al. (2021); Gao et al. (2023b); Liu et al. (2024). Ever since the advent of Large Language Models (LLMs), such as GPT-3 Brown et al. (2020) or LLAMA-2 Touvron et al. (2023), they are the de-facto choice for answer generation in RAG, due to their unprecedented progress in text generation and understanding Brown et al. (2020); Li et al. (2024); Chang et al. (2024); Guo et al. (2023). RAG grounds the LLM-generated answer, thereby avoiding previously observed factual hallucination Maynez et al. (2020); Raunak et al. (2021) and outdated knowledge Cao et al. (2021); He et al. (2023) in LLMs.
@@
A challenging issue in RAG is to provide robust and reliable LLM-generated answers. The answer generation stage is dependent on the first-stage information retrieval system. The retrieval system poses challenges in accurately retrieving relevant information when evaluated either on zero-shot domains Thakur et al. (2021) or low-resource languages Zhang et al. (2023). The incorrect or non-relevant information contained in retrieved passages can frequently mislead the LLM to hallucinate Adlakha et al. (2024); Chen et al. (2024); Shi et al. (2023); Yoran et al. (2024); Yu et al. (2023). Prior work lacks a comprehensive evaluation of LLM reasoning capabilities in multiple languages. As a result, it remains unclear to which extent LLMs hallucinate across both high- or low-resource languages.
To facilitate research in this direction, we present NoMIRACL, a large multilingual human-annotated dataset containing over 56,000 (including both non-relevant and relevant samples) to evaluate LLM robustness against errors in first-stage external information, i.e., retrieved passages, across 18 typologically diverse languages. To construct the dataset, we hired 31 native speakers as human annotators Zhang et al. (2023). NoMIRACL contains two subsets, non-relevant and relevant. The non-relevant subset contains all queries with no known answers, i.e., all top-$k$ retrieved passages manually judged as non-relevant. Conversely, the relevant subset contains queries with known answers, i.e., at least one of the top-$k$ passages is manually judged as relevant.
To better understand the LLM robustness in NoMIRACL, we conduct experiments with several existing powerful and multilingual-focused LLMs (e.g., GPT-4, Mistral, LLAMA-3). We conduct our experiments using the top-$k$ oracle passages retrieved using a hybrid retrieval system from a language-specific Wikipedia corpus Zhang et al. (2023). We use a zero-shot “vanilla” prompt template for prompting all LLMs. Our key findings are: First, LLMs such as LLAMA-2, Aya-101, and Orca-2 observe a surprisingly high 88% hallucination rate on the non-relevant subset. Second, the Mistral and LLAMA-3 series of models hallucinate less but perform worse on the relevant subset. Overall, GPT-4 is found to provide the optimal performance tradeoff across both subsets.
To understand our experimental findings better, we conduct an empirical analysis on NoMIRACL (en) to analyze the blind spots in a subset of LLMs. We observe LLAMA-2-7B and 13B interestingly repeat the query and prompt instructions on average by at least 25%. Mistral-7B always provides a rationale in their output generation by over 88%. In addition, we conduct different prompting techniques and observe that supervised fine-tuning LLMs on the NoMIRACL development set can be tricky.
@@ 2 Background and Problem Identification
RAG Background. Retrieval-augmented generation (RAG) (Lewis et al., 2020; Guu et al., 2020) involves a two-stage inference pipeline. In the first stage, given the retrieval system and the user query, the retrieval system provides the subset of top-$k$ passages retrieved from an external data corpus $C$. For the next stage, the user query with the top-$k$ retrieved passages is provided to the LLM, which generates an output summarizing the answer for the query and citing the relevant passages.
@@ 2.1 Robustness Evaluation
Definitions. NoMIRACL contains two subsets, we denote them as either non-relevant ($\mathrm{F}$) or relevant ($\mathrm{T}$). The non-relevant subset contains queries with no-known answers, i.e., all top-$k$ passages are non-relevant, while the relevant subset contains queries with known answers, i.e., at least one of the top-$k$ passages is relevant. Similarly, we denote the LLM prediction as either positive ($\mathrm{P}$) indicating the model finds the passage relevant to answer the query and similarly negative ($\mathrm{N}$) denotes the model does not find any passage relevant (i.e., containing the answer) for the query.
@@ 3.1 Data Construction Procedure
NoMIRACL is constructed using the same procedure utilized to develop MIRACL Zhang et al. (2023). The data construction occurs in two stages, following Zhang et al. (2023); Clark et al. (2020). In the first stage, the annotator (a native language speaker) writes a well-formed query for each individual prompt text. Each prompt is a short text snippet containing the first 100 words from a language-specific Wikipedia corpus. Next, for each human-generated query, top-$k$ passages are retrieved from the corpus using a hybrid multilingual retrieval system (more details in §4.1). In the second stage, annotators assess the binary relevance judgment of the top-$k$ query–passage pairs, either relevant (relevance = $1$) or non-relevant (relevance = 00). For additional details in data construction, such as quality control, we would like to refer the reader to Appendix C.
@@
Relevant Subset. Queries with known answers, i.e. at least one of top-$k$ retrieved passages marked by the annotator as relevant to provide sufficient information to answer the query. We construct the NoMIRACL relevant subset with queries with at least one relevant passage, i.e., all query-passage pairs have been judged as either relevant with a relevance of $1$ or non-relevant with relevance of 00.
@@ 3.3 Languages Covered and Dataset Usage
NoMIRACL covers 18 diverse typological languages Zhang et al. (2023). [^3] NoMIRACL covers 10 families (from Niger-Congo to Indo-European) and 11 scripts (from Latin to Devanagari) covering diversity from the perspective of linguistic characteristics. The languages along with their ISO codes are: Arabic (ar), Bengali (bn), German (de), English (en), Spanish (es), Persian (fa), Finnish (fi), French (fr), Hindi (hi), Indonesian (id), Japanese (ja), Korean (ko), Russian (ru), Swahili (sw), Thai (th), Yoruba (yo), Chinese (zh).
From Table 1, we observe an uneven amount of queries present in each language. To avoid this non-uniformity and budget constraints (see subsection 4.3), in our experiments, we limit the maximum number of 250 queries for each language and subset (if available) in NoMIRACL.
@@ 4.1 Evaluation Setup and Metrics
In NoMIRACL, we assess LLM relevance as either hallucination or error, using an input query, a vanilla prompting technique, and top-$k$ (oracle) retrieved and relevance judged passages.
@@
Evaluation Objective. In our work, we evaluate LLM relevance as a response string $y$ in a binary classification setup. Following prior evaluation strategies in Adlakha et al. (2024); Yu et al. (2023), we use the input query $q_{i}$, a vanilla prompt template $Q$, and a set of top-$k$ annotated passages $P_{k}$. We prompt the LLM to evaluate if $q_{i}$ can be answered using any passage in $P_{k}$. The LLM generates an answer output containing either $y$ = “I don’t know” as negative ($\mathrm{N}$) or “Yes, answer is present” as positive ($\mathrm{P}$). The output is tagged “Invalid” if it does not fall in either one of the above. Recall from (§2.1), we calculate the hallucination rate (in %) = $\mathrm{FP}/(\mathrm{FP}+\mathrm{TN})$, which measures the error in rejecting information from non-relevant passages and the error rate (in %) = $\mathrm{FN}/(\mathrm{FN}+\mathrm{TP}$), which measures the error in identifying relevant passages amongst noisy ones.
@@ 4.2 Evaluation Models
We evaluate eleven state-of-the-art LLMs with a strong focus of multilingual instruction capabilities, including both open and closed-sourced. All model checkpoints can be found in Appendix E.
(1) OpenAI: We include three closed-book LLM variants: GPT-3.5-turbo, GPT-4, and GPT-4o OpenAI (2023) using the Azure OpenAI service. (2) Mistral: We include two variants: (i) Mistral-7B-Instruct-v0.3, the latest 7B instruction-tuned parameter model Jiang et al. (2023) and (ii) Mixtral-8x7B-v0.1, a sparse Mixture-of-Expert (MoE) model Jiang et al. (2024). (3) Orca-2: In the Orca-2 series Mitra et al. (2023), we include both Orca-2-7B and Orca-2-13B. (4) Aya: Aya-101 Üstün et al. (2024) is a recently introduced multilingual LLM containing 13B parameters and trained with 101 languages and Aya-23-35B finetuned across 23 languages Aryabumi et al. (2024). (5) LLAMA-2: In the LLAMA-2 series Touvron et al. (2023), we include three chat variants: Llama-2-7b-chat-hf, Llama-2-13b-chat-hf, and Llama-2-70b-chat-hf instruction tuned chat models. (6) LLAMA-3: Following up on the LLAMA-2 series, we include both the instruction tuned models Dubey et al. (2024): Meta-Llama-3-8B-Instruct and Meta-Llama-3-70B-Instruct.
@@ 4.3 Experimental Settings
We execute the generation of GPT-4o, GPT-4, and GPT-3.5-turbo, using the OpenAI service (API version 2023-05-15) deployed on Microsoft Azure [^4] We compared Azure API with OpenAI API across four languages in NoMIRACL and observed no noticeable difference between different GPT-4 API version providers. and LLAMA-3 series using the AnyScale API service. We maintain a maximum input sequence length of 4096 tokens for a fair evaluation amongst all models. We set a low-temperature score = 0.1 for a deterministic output, and a top-p sampling ratio = 0.95. We output a maximum of $50$ tokens.
{mdframed}
[backgroundcolor=gray!5] I will give you a question and several contexts containing information about the question. Read the contexts carefully. If any of the contexts answers the question, respond as either ‘‘Yes, answer is present’’ or ‘‘I don’t know’’:
@@
Figure 5: Hallucination rate (in %) = $\mathrm{FP}/(\mathrm{FP}+\mathrm{TN}$) on the non-relevant subset ($\mathrm{F}$) in NoMIRACL test split. The non-relevant subset contains queries with no known answers, i.e., all top-$k$ (where $k=10$) passages are judged by a human annotator as non-relevant. A majority of LLMs (except Mistral) hallucinate on the non-relevant subset. Lower the hallucination rate is better. The best model in each category is plotted (see Figure 8 for all models).
@@
Vanilla Prompting. The choice of prompt significantly influences the performance and LLMs have been shown brittle to prompting variations, training examples, or long context setups Liu et al. (2024). In our work, we evaluate all baselines using a zero-shot monolingual listwise prompting strategy. We construct a vanilla prompt template using all top-$k$ (oracle) passages (available in NoMIRACL) as a list of contexts along with the input query, both in the same language. We evaluate all LLMs zero-shot, as we cannot fit few-shot exemplars due to insufficient context length (maximum of 4096 sequence length). Our template provides a short description in English describing the task (Ahuja et al., 2023). Our vanilla prompt template used in our experiments is shown in Figure 4.
Reducing Costs. Running GPT-4 is expensive and LLAMA-3-70B is rather slow at inference. For long contexts and low-resource languages, the costs can even multiply. To limit, we did not exceed our prompt above 4096 tokens. To effectively fit all $k=10$ passages within the vanilla zero-shot prompt, we truncate each passage to use the first 375 tokens. Next, due to budget constraints, we keep our evaluation to a maximum of 250 randomly sampled queries for all languages in both NoMIRACL relevant and non-relevant split. We end up providing $\approx$ 20K API calls producing an expense of \$1,474 (in USD) including miscellaneous costs.